A 25.00 mL sample of .0056 M triethylamine (N(C2H5)3 pKb= 3.25) is titrated with standardized .0056 M HCL solution. What is the pH of the solution after 12.5 mL of HCl has been added? (Hint- half way to equivalence point)n
Answer #1

Addition of 12.5 mL of HCI: Number of moles of (CH3)2N = M*V = 0.0056 M* 25.00 mL = 0.14 mmol Number of moles of HCl = M*V =